An aquarium of length L, width (front to back) W, and depth D is filled to the top with liquid of density ?.

Part A) Evaluate the forces on the front window for a 95-cm-long, 25-cm-wide, 40-cm-deep aquarium filled with water.

Part B)Evaluate the forces on the bottom for a 95-cm-long, 25-cm-wide,40-cm-deep aquarium filled with water.
Answer #1

Part A) resultant force on front window

   F=(P+\rho gh/2)*Area

F=(101325+1000*9.81*.2)*0.2375=24530 N=24.53 kN

Part B) force on the bottom

F=(P+\rho gh)*Area

   F=(101325+1000*9.81*.4)*0.2375=24996.6N=24.9966kN